OpenStudy (anonymous):

|sin(x)-3|=2 whats x?

OpenStudy (anonymous):

the only way for that to happen is if \(\sin(x)=1\)

OpenStudy (anonymous):

the full questions is : find the value of x in the interval 0<x<360 that satisfies the equation |sin(x)-3|=2

OpenStudy (anonymous):

same answer only true if \(\sin(x)=1\)

OpenStudy (anonymous):

so how do you solve it ?

OpenStudy (anonymous):